Solve for y:
xy = 1 + 3y

Answer by Alan3354(69443)

Do you have any ideas?
--------------
Step 1 - get all the y terms on one side of the eqn.
xy = 1 + 3y
xy-3y = 1
Step 2 - factor out the y if possible.
y(x-3) = 1
y = 1/(x-3)